OpenStudy (anonymous):

Joe and his friends are playing a game w/ 2 dice. Joe must roll a sum of 7 to get points. Jay must roll a sum of 5. Jane must roll a sum of 2, and Jamie a sum of 10. Which friend is more likely to get the most points?

OpenStudy (anonymous):

@Kitt020912

OpenStudy (tkhunny):

Standard Dice: p(2) = p(12) = 1/36 p(3) = p(11) = 2/36 Can you complete the table?

OpenStudy (anonymous):

uhm, let me try

OpenStudy (anonymous):

Chinchilla, everyone else is beating me to the answers I'm sorry :l

OpenStudy (anonymous):

It's ok:)

OpenStudy (anonymous):

Francais, oui?

OpenStudy (anonymous):

Could you explain more thoroughly what you want me to do, Tk?

OpenStudy (tkhunny):

Who has the highest probability? p(2) = p(12) = 1/36 p(3) = p(11) = 2/36 p(4) = p(10) = 3/36 p(5) = p(9) = 4/36 Can you finsh the table, now?

OpenStudy (anonymous):

Lol, vous aimez parler français aussi bien que vous n'avez pas? :D

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

Je n'ai toujours pas l'obtenir

OpenStudy (anonymous):

Et vous?

OpenStudy (tkhunny):

En halua puhua ranskaa. I am not yet asking you to get it. I am asking you to complete the table. If we have the table, we can solve the problem.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

You speak finnish? :0

OpenStudy (anonymous):

anyway.... let me attempt it...again...

OpenStudy (anonymous):

p(6) = p(8) = 3/36 like that?

OpenStudy (tkhunny):

Osaan hiukan. p(6) = p(8) = 5/36 p(7) = 6/36 Joe must roll a sum of 7 to get points. Jay must roll a sum of 5. Jane must roll a sum of 2, and Jamie a sum of 10. p(Joe gets a point on a single roll) = 6/36 p(Jay gets a point on a single roll) = 4/36 p(Jane gets a point on a single roll) = 1/36 p(Jamie gets a point o na single roll) = 3/36 Are we yet seeing it?

OpenStudy (anonymous):

er, let me read all thet^ one moment

OpenStudy (anonymous):

where are you getting the numbers from?

OpenStudy (anonymous):

o right, but y are u subtracting one?

OpenStudy (tkhunny):

Draw a 6x6 grid. At the top of each column, write the values 1, 2, 3, 4, 5, 6 At the left of each row, write the values 1, 2, 3, 4, 5,6 Fill in the grid with the values obtained by adding the corresponding column header and row header. See if your grid looks like a good representation of what you would get by rolling two standard dice. Count the 2s in the grid. Count the 3s in the grid. Count the 4s in the grid. ... Count the 11s in the grid. Count the 12s in the grid.

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

@Kitt020912

OpenStudy (anonymous):

s'il vous plaît aidez-moi

OpenStudy (anonymous):

hmm. just a sec alright?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

@Firejay5

OpenStudy (anonymous):

[kitt: webwhiteboard <--]

OpenStudy (anonymous):

Help please.

OpenStudy (anonymous):

mitt one moment..

OpenStudy (firejay5):

Joe will have the most possibilities

OpenStudy (anonymous):

How so?

OpenStudy (firejay5):

Well you have to think of it as dice(die)

OpenStudy (anonymous):

Okay.

OpenStudy (firejay5):

6 sides of a dice

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

And then?

OpenStudy (firejay5):

Follow this: 1st dice 2nd dice 1 1 2 2 3 3 4 4 5 5 6 6

OpenStudy (anonymous):

Hmmm. Okay.

OpenStudy (anonymous):

So Joe would be 6 6 ?

OpenStudy (firejay5):

Joe would have the most possibilities

OpenStudy (anonymous):

Alright.

OpenStudy (firejay5):

Adding number that add to 7: 1 and 6, 5 + 2, and 3 and 4

OpenStudy (anonymous):

Ok, thanks!

OpenStudy (anonymous):

I have some more.

OpenStudy (firejay5):

Joe = 3 Jay = 2 Jane = 1 Jamie = 2

OpenStudy (anonymous):

Oooh. Okay.

OpenStudy (firejay5):

do you want to know why Jay, Jane, and Jamie aren't it

OpenStudy (anonymous):

u have to see how many numbers on a die would add up to each certain number,yes?

OpenStudy (firejay5):

yes probability is basically like adding up, but not always

OpenStudy (anonymous):

Okay. Why are the others not it?

OpenStudy (firejay5):

Jay is sum of 5 right

OpenStudy (anonymous):

Yes.

OpenStudy (firejay5):

well there's a 2 probability of getting a 5: 3 + 2 and 4 + 1

OpenStudy (anonymous):

Ohhhhhh. OKAY!

OpenStudy (anonymous):

I think Im understanding.

OpenStudy (firejay5):

Jane is a sum of 2: 1 + 1 Jamie is a sum of 10: 5 + 5 and 6 + 4

OpenStudy (firejay5):

How old are you and what grade are you in

OpenStudy (anonymous):

Ooooh. How many combinations of the certain amount of points.

OpenStudy (anonymous):

Why and why?

OpenStudy (anonymous):

Anyway, thanks. I'm off to post some more questions. :)

OpenStudy (firejay5):

just asking sorry if it's too personal

OpenStudy (anonymous):

Oh. Just because?

OpenStudy (anonymous):

Well, let me finish posting this question first.

OpenStudy (firejay5):

just to see what grade you are learning this in

OpenStudy (firejay5):

about 5th - 8th grade